OpenStudy (anonymous):

i need help rewriting this equation with only sin x and cos x : sin 3x - cos x

OpenStudy (phi):

sin(a+b)= sin(a)cos(b)+sin(b)cos(a) cos(a+b)= cos(a)cos(b)-sin(a)sin(b) let a= 2x and b= x then do it again with a=x and b=x to get rid of the 2x